OpenStudy (anonymous):

use combinatorics to find the 5th term of (2+z)^8

OpenStudy (anonymous):

= 8C4 2^4 *z ^4 = 1120 z ^4

OpenStudy (anonymous):

\[\left(\begin{matrix}8 \\ 4\end{matrix}\right)\] nCr = \[n! \over (n - r)! r!\] 8C4 = \[8! \over (8 - 4)! 4!\] \[8 * 7 * 6 * 5 \over4 * 3 * 2 * 1\] 8C4 = 70

OpenStudy (anonymous):

That's your coefficient. Just plug in the numbers now :)

sam (.sam.):

or you can use my guide here http://openstudy.com/study?login#/updates/4fb7bd6fe4b05565342d33a0

OpenStudy (anonymous):

70(2)^4(z)^4 = 1120z